OpenStudy (anonymous):

how do i do this? 2+x/4=-3

OpenStudy (anonymous):

multiply all terms by 4 and then isolate x

OpenStudy (anonymous):

8+x=-12

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i do this? 3/5x-1=1

OpenStudy (anonymous):

multiply everything by 5

OpenStudy (anonymous):

3x-5=5

OpenStudy (anonymous):

it will be x=2/3. right?

OpenStudy (anonymous):

x=10/3

OpenStudy (anonymous):

thank you